OpenStudy (123ab456c):

The ratio of two numbers is 2:3, and their sum is 60. Find the two numbers. A) 10 and 50 B) 12 and 48 C) 20 and 30 D) 24 and 36

OpenStudy (123ab456c):

just EXPLAIN !!

jhonyy9 (jhonyy9):

let these two number x and y ratio : x/y = 2/3 x+y=60 solve this system for x and y

jhonyy9 (jhonyy9):

can you solve it please ?

OpenStudy (123ab456c):

2/3

OpenStudy (123ab456c):

@jhonyy9

jhonyy9 (jhonyy9):

2/3 is the ratio between x and y

jhonyy9 (jhonyy9):

there is a system what you just need to solve it for x and y - ok. ?

jhonyy9 (jhonyy9):

from x+y=36 => x= 36-y so than this x=36-y substitute in place of x inside the first equation - ok. ?

jhonyy9 (jhonyy9):

sorry my mistaqe not 36 there is 60 - ok. ?

jhonyy9 (jhonyy9):

@ghgbf do you can solve it for medal please ?